OpenStudy (anonymous):

Solve the integral: 1 / (e^x + 1) Explain thoroughly the algebraic step that is necessary.

OpenStudy (amistre64):

hmm....

OpenStudy (amistre64):

\[\int\frac{1}{e^x+1}dx\] if we had an e^x up top we could ln it

OpenStudy (amistre64):

do we multiply by a conjugate?

OpenStudy (amistre64):

add zero to the top?

OpenStudy (amistre64):

add zero is what id try

OpenStudy (amistre64):

\[\int\frac{1+(e^x-e^x)}{e^x+1}dx\] \[\int\frac{1+e^x-e^x}{e^x+1}dx\] \[\int\frac{1+e^x}{e^x+1}-\frac{e^x}{e^x+1}dx\] \[\int\ 1-\frac{e^x}{e^x+1}dx\] \[x - ln(e^x+1)+C\]

OpenStudy (anonymous):

Nicely done, thanks.